OpenStudy (anonymous):

Please help me! WRITTEN RESPONSE--- Bryan has to carry 24 books to a new shelf. If he carries one book at a time, it will take him 24 trips. If he carries 3 books at a time, how many trips will it take? Does this relationship represent direct or inverse variation? Using the appropriate formula, find k. Show your work and answer the original question

OpenStudy (anonymous):

If B carries 3 books at a time in 1st trip he puts 3 in the shelf, in 2nd 6, in 3rd 9... Do you understand how it works?

OpenStudy (anonymous):

yes , but what about the questions at the bottom. thats what I need help with.

OpenStudy (anonymous):

Well, he will go until all the books are in the shelf. If he carries 1 book at a time he goes 24 times, if he carries 3 books, he goes x times. if he carries more books at a time he goes less times. (inverse).

OpenStudy (anonymous):

so what would be the final formula or equation to this question?

OpenStudy (anonymous):

3*x=24.

OpenStudy (anonymous):

so inverse , formula of finding k is 3*x=24 , and original question anwser is 3

OpenStudy (anonymous):

sorry original question anwser 8

OpenStudy (anonymous):

Number of trips =number of books / books in a trip ( y = 24/x ) that´s inverse, and the answer is 8.

OpenStudy (anonymous):

so thats the question anwsers at the bottom mainly

OpenStudy (anonymous):

like what would I write exactly as the anwsers in order at the bottom. srry

OpenStudy (anonymous):

a) 24/3 = 8 b)Inverse.. c) y = (1/x)*24. So k = 1/x. d) y is number of trips, 1/x is k and 24 are the total books B must carrie.

OpenStudy (anonymous):

thank you so much

OpenStudy (anonymous):

Glad to help